A test to determine whether a certain antibody is present is 99.2 ​% effective. This means that the test will accurately come back negative if the antibody is not present​ (in the test​ subject) 99.2 ​% of the time. The probability of a test coming back positive when the antibody is not present​ (a false​ positive) is 0.008 . Suppose the test is given to four randomly selected people who do not have the antibody. ​(a) What is the probability that the test comes back negative for all four ​people? ​(b) What is the probability that the test comes back positive for at least one of the four ​people?
Answer #1

(a)

BinomialDistribution

n= 4

p = Probability that the test comes back negative = 0.992

q = 1 - p = 0.008

PlX = 4) = 0.00800.9924-1 × 1 × 0.9684 0.9684

(b)

P(test comes back positive for at least 1)= 1 - P(Test comes back negative for all four people)

                                                            = 1 - 0.9684 = 0.0316